I would like to know the answers to these questions:
m to the third, n squared, both inside a radical sign
50, inside a radical sign
8 in a radical sign, multiplying a 7 in a radical sign

Answer by checkley75(3666)

sqrt(m^3n^2)=sqrt(m^2*m*n^2)=mnsqrt(m)
sqrt50=sqrt(25*2)=5sqrt(2)
sqrt8*sqrt7=sqrt(4*2)*sqrt(7)=2sqrt(2)*sqrt(7)